OpenStudy (anonymous):

What is the aswer for. y-5 ----- = 3 3 Please help i dont want to get kicked off the team for bien an eligilble..

OpenStudy (ashley1nonly):

is that (y-5)/3 =3

OpenStudy (ashley1nonly):

??????????

OpenStudy (ashley1nonly):

if so first you will multiply both sides by 3 to remove the 3 from the bottom equation

OpenStudy (ashley1nonly):

now you have 3(y-5)/3 = 3(3) do that and get y-5=9 add 5 to both sides y=14 plug 14 in four y 14-5=9 9/3=3 y=14

OpenStudy (ashley1nonly):

medal